# Flow-Inspired Lightweight Multi-Robot Real-Time Scheduling Planner

### v1

@@ VI Experiments
This section provides an experimental analysis of our method. Our experiments are conducted on a PC with an Inter Core-i7 12700 CPU under 32GB of RAM. We compared the following methods to verify the performance of the proposed method:
(1) Astar: Each robot uses the Astar algorithm to calculate the path from the starting point to the goal point.
(2) Greedy: Find all paths from the upper boundary node of the starting cell to the lower boundary node of the goal cell. Then, use a greedy approach to evenly distribute robots across each path according to the path length and capacity.
(3) RunCost: The method is almost the same as FRSP. But the objective function only has running cost. The coefficients are set to $k_{1}=0,k_{2}=0,k_{3}=1$.
(4) FRSP: Our scheduling planner. The coefficients are set to $k_{1}=1,k_{2}=0.5,k_{3}=0.5$. The MIQP is solved by Gurobi [18].
The maps are categorized into two types: random forest and maze, with 5 maps generated for each type. For each map, we conducted experiments with varying numbers of robots, specifically: 10, 50, 100, 150, 200, 250, 300, 350, 400, 450, 500. The initial and goal positions of the robots are uniformly arranged in a matrix queue format. The parameters for the experiments are set as follows: $r_{\min}=0.4$m, $\alpha=2$, $h=0.01s$, $\mathbf{v}_{\max}=3$m/s, $N_{B}=4$. Our local collision avoidance algorithm utilizes the Reciprocal Velocity Obstacles (RVO) [3] method, chosen for two primary reasons. Firstly, RVO demonstrates reliable performance in multi-robot collision avoidance, reducing the likelihood of deadlocks. Secondly, it offers high computational efficiency, maintaining acceptable computation times even with a large number of robots. The frequency of collision avoidance algorithm is 100 Hz and the scheduling is 1 Hz.
@@
Figure 6: Time results w.r.t. number of drones
Fig.6 shows the time results with respect to the number of robots in forest and maze maps. For every number of robots $N$, mean values are reported for each type of 5 maps. From the figures, we can make the following observations: (1) When $N$ is small, there is not much difference between the time of four methods. This is because when the number is small, congestion rarely occurs, so the time results of the methods are similar. (2) As $N$ increase, the time results increase. The gap between time results of FRSP and other methods is gradually increasing. This is because as $N$ increases, the congestion starts to worsen, and the robots will waste more time waiting. FRSP can plan new paths for the waiting robots, thus reducing the time it takes to cross. The larger $N$ is, the better FRSP works. (3) FRSP performs better on forest maps than on maze maps. This is because the forest map is more complex than the maze map, and there are more forks in the network structure. This shows that FRSP can adapt to complex map structures, and the more complex the map structure, the better the effect.
The average improvemnt
Astar RunCost Greedy
FRSP Forest 8.66% 7.52% 10.77%
Maze 6.58% 4.68% 7.54%
@@
Table.I shows the average improvement of FRSP. The improvement ratio of FRSP in forest map compared with the other three methods are 8.66%, 7.52%, 10.77%. The improvement ratio of FRSP in maze map are 6.58%, 4.68%, 7.54%. It can be seen that the Greedy method is not as good as RunCost and Astar. This is because both RunCost and the Astar choose the shortest distance path, and they will wait when faced with congestion. But Greedy assigns robots to each path according to the greedy idea. Although congestion was reduced, time was wasted on further travel, suggesting that sometimes waiting may be a good option. Table.II shows the computation time with respect to the number of robots $N$. It can be seen that: (1) Path set search and path node selection take up most of the computation time, while local position allocation takes very little computation time. (2) When $N$ is 10, the computation time of FRSP is about 0.06 seconds. When $N$ is 500, the computation time of FRSP is about 0.9 seconds. This shows that our FRSP is computationally efficient and meets the real-time requirements.
@@ VI-A Real World Flight Test
We conduct flight tests with ten Dij Tello drones within a forest environment. The drones are constrained to a space of $15m\times 5m\times 2m$ to prevent them from completely circumventing obstacle areas. All computations are performed on a single laptop. The supplementary video demonstrates the complete execution of the test, with no collisions or deadlocks occurring throughout the flight.
@@ VII Conclusion
In this paper, we proposed a flow inspired scheduling planner to optimize the traversal of multi-robot systems through obstacle-rich environments. Simulation and experimental results demonstrate the effectiveness of our scheduling planner.

### v2

@@ VI Experiments
((a)) $t=1.4s$
((b)) $t=4.8s$
((c)) $t=6.2s$
((d)) $t=10.4s$
Figure 7: The scheduling process of 500 robots in the forest map.
In this section, we conduct two sets of experiments to explore the performance of our method in multi-robot planning. All experiments are carried out on a PC equipped with an Intel Core-i7 12700 CPU and 32GB of RAM.
The number of robots in the first set of experiments ranges from 10 to 100. (10, 20, 30, 40, 50, 60, 70, 80, 90, 100) For each number of robots, we conduct experiments on two different map types: random forest and maze. Five different instances are generated for each map type. In this set of experiments, the compared algorithms are as follows:
  • Astar: Each robot uses the Astar algorithm to calculate the path from the starting point to the goal point.
  • ECBS (Enhanced Conflict Based Search) [17]
  • P-SIPP (Prioritized Safe Interval Path Planning) [18]
We use FRSP to denote our algorithm. The coefficients of FRSP are set to $k_{1}=1,k_{2}=0.5,k_{3}=0.5$. The MIQP is solved by Gurobi [19].
In the second set of experiments, we explore the impact of a large number of robots on the algorithms. The number of robots is set to 100, 150, 200, 250, 300, 350, 400, 450, and 500. We also conduct experiments on two types of maps: forest and maze. However, during the experiment, it is found that when the number of robots is higher than 100, the success rates of the ECBS and P-SIPP algorithms are extremely low, making effective comparison impossible. Therefore, in this set of experiments, the following three methods are selected for comparison:
  • Astar: Each robot uses the Astar algorithm to calculate the path from the starting point to the goal point.
  • Greedy: Find all paths from the upper boundary node of the starting cell to the lower boundary node of the goal cell. Then, use a greedy approach to evenly distribute robots across each path according to the path length and capacity.
  • RunCost: The method is almost the same as FRSP. But the objective function only has running cost. The coefficients are set to $k_{1}=0,k_{2}=0,k_{3}=1$.
Two sets of experiments are set up with the aim of demonstrating that our algorithm not only outperforms the commonly used ECBS and P-SIPP algorithms within the number range of 10 to 100. In the large scale experiments (100 to 500) where both ECBS and P-SIPP are unable to cope, our algorithm can still handle and exhibit good performance.
The parameters for the experiments are set as follows: $r_{\min}=0.4$m, $\alpha=2$, $h=0.01s$, $\mathbf{v}_{\max}=3$m/s, $N_{B}=4$. The initial and goal positions of the robots are uniformly arranged in a matrix queue format. Our local collision avoidance algorithm utilizes the Reciprocal Velocity Obstacles (RVO) [1] method, chosen for two primary reasons. Firstly, RVO demonstrates reliable performance in multi-robot collision avoidance, reducing the likelihood of deadlocks. Secondly, it offers high computational efficiency, maintaining acceptable computation times even with a large number of robots. The frequency of collision avoidance algorithm is 100 Hz and the scheduling is 1 Hz.
@@
Figure 8: Time results w.r.t. number of robots (10 to 100)
((a)) Forest maps
((b)) Maze maps
Figure 9: Time results w.r.t. number of robots (100 to 500)
Fig.8 shows the time results for the number of robots ranging from 10 to 100 in the forest and maze maps, while Fig.9 shows the time results for the number of robots ranging from 100 to 500 in the forest and maze maps. For every number of robots $N$, mean values are reported for each type of 5 maps. We can observe that: (1) The time result of FRSP is the best, while that of P-SIPP is the worst. The time results of ECBS and Astar are similar, with ECBS being slightly inferior to Astar. This indicates that the FRSP can better handle the congestion caused by the increase in the number of robots. (2) The Astar algorithm searches for the shortest path to the destination for each robot. ECBS and P-SIPP, on the other hand, search for discrete, conflict-free paths for the robots. However, these paths do not take into account the impact of congestion. As a result, as the value of $N$ increases, the time efficiency decreases. (3) FRSP still performs well in the range of 100 to 500. The performance of Greedy algorithm is the worst. (4) When $N$ is small, there is not much difference between the time of four methods. This is because when the number is small, congestion rarely occurs, so the time results of the methods are similar. (5) As $N$ increase, the time results increase. The gap between time results of FRSP and other methods is gradually increasing. This is because as $N$ increases, the congestion starts to worsen, and the robots will waste more time waiting. FRSP can plan new paths for the waiting robots, thus reducing the time it takes to cross. The larger $N$ is, the better FRSP works. (6) FRSP performs better on forest maps than on maze maps. This is because the forest map is more complex than the maze map, and there are more forks in the network structure. This shows that FRSP can adapt to complex map structures, and the more complex the map structure, the better the effect.
The average improvement
Range from 10 to 100 Range from 100 to 500
Astar ECBS P-SIPP Astar Greedy RunCost
FRSP Forest 9.04% 11.37% 14.78% 8.66% 7.52% 10.77%
Maze 7.07% 8.13% 11.57% 6.58% 4.68% 7.54%
@@
Table.I shows the average improvement of FRSP. It can be seen that: (1) the Greedy method is not as good as RunCost and Astar. This is because both RunCost and the Astar choose the shortest distance path, and they will wait when faced with congestion. But Greedy assigns robots to each path according to the greedy idea. Although congestion was reduced, time was wasted on further travel, suggesting that sometimes waiting may be a good option. (2) In the scenario of multi-robots moving from one side to the other side of an obstacle-filled area, the ECBS and P-SIPP algorithms perform even worse than the Astar algorithm. This is in line with expectations because in the algorithm design of ECBS and P-SIPP, the initial and goal positions of the robots are scattered, and these scattered positions are less likely to cause congestion. This indicates that the commonly used ECBS and P-SIPP are highly unsuitable for addressing congestion situations.
Table.II shows the computation time with respect to the number of robots $N$. It can be seen that: (1) Path set search and path node selection take up most of the computation time, while local position allocation takes very little computation time. (2) When $N$ is 10, the computation time of FRSP is about 0.06 seconds. When $N$ is 500, the computation time of FRSP is about 0.9 seconds. This shows that our FRSP is computationally efficient and meets the real-time requirements.
@@ VI-A Real World Flight Test
Figure 10: Real flight test scene.
Figure 11: The flight process of 10 drones.
We conduct flight tests with ten Dij Tello drones within a forest environment. The drones are constrained to a space of $15m\times 5m\times 2m$ to prevent them from completely circumventing obstacle areas. We use OptiTrack to provide high-precision positioning information for drones. Fig.10 shows our real flight test scene where we placed 5 boxes as obstacles in the scene. The parameter settings of the real flight scene are as follows: $r_{\min}=0.5$m, $\alpha=1$, $h=0.01s$, $\mathbf{v}_{\max}=1$m/s, $N_{B}=1$. The frequency of collision avoidance algorithm is 100 Hz and the scheduling is 1 Hz. Fig.11 shows the flight process of 10 drones. All calculations are performed on a single laptop and no collisions or deadlocks occurred during the entire flight. The detailed process can be seen in the supplementary video.
@@ VII Conclusion
In this paper, we proposed a flow inspired scheduling planner to optimize the traversal of multi-robot systems through obstacle-rich environments. Simulation and experimental results demonstrate the effectiveness of our scheduling planner. In the future work, we will consider how to perform map decomposition and network construction under obstacles of arbitrary shapes. Try different collision avoidance algorithms to enhance the applicability of our scheduling planner. And in the real world experiment, we will try to conduct more drones tests and try more complex environments.